OpenStudy (anonymous):

for the simple harmonic motion equation d=9 cos(pi/2t) what is the period?

OpenStudy (anonymous):

is it 1?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

period of \(\cos(bx)\) is \[\frac{2\pi}{b}\]

OpenStudy (misty1212):

so no, it is not one

OpenStudy (anonymous):

so 2pi/b?

OpenStudy (misty1212):

lol no dear, you need a number out of this in your case \(b=\frac{\pi}{2}\) so divide \(2\pi\) by \(\frac{\pi}{2}\)

OpenStudy (anonymous):

4?

OpenStudy (misty1212):

yes four !

OpenStudy (anonymous):

So then 4 would be the period?

OpenStudy (misty1212):

yes 4 would be the period

OpenStudy (anonymous):

Thank you!:)

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]